# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ Appendix F Relative Increase R
$\mathbb{E}_{x^{\prime},{y}}\mathbb{E}_{\mathcal{T}}[\|y-f_{n}(x^{\prime})\|^{2}_{2}]$ $=\underbrace{\mathbb{E}_{x^{\prime},{y}}[\|y-\bar{f_{n}}(x^{\prime})\|^{2}_{2}]}_{\text{Bias}^{2}}+\underbrace{\mathbb{E}_{x^{\prime},\mathcal{T}}[\|\bar{f}_{n}(x^{\prime})-f_{n}(x^{\prime})\|^{2}_{2}]}_{\text{Variance}}$ (30)
$=\mathbb{E}_{x^{\prime},{y}}[\|y-\bar{f}(x^{\prime})\|^{2}_{2}]+\frac{1}{n}\mathbb{E}_{x^{\prime},\mathcal{T}}[\|\bar{f}(x^{\prime})-f(x^{\prime},\mathcal{T})\|^{2}_{2}]$ (31)
Inequality 11 holds due to that $\text{Var}(\frac{1}{n}\sum_{i=1}^{n}f_{i}(\cdot;\mathcal{T}_{i}))=\frac{1}{n^{2}}\text{Var}(\sum_{i=1}^{n}f_{i}(\cdot;\mathcal{T}_{i}))$ and since each $f_{i}$ is trained independently, $\frac{1}{n^{2}}\text{Var}(\sum_{i=1}^{n}f_{i}(\cdot;\mathcal{T}_{i}))=\frac{1}{n^{2}}\sum_{i=1}^{n}\text{Var}(f(\cdot;\mathcal{T}))=\frac{1}{n}\text{Var}(f(\cdot;\mathcal{T}))$. ∎
Appendix I linear regression to visualize the correlation
@@
(a) ResNet50: transfer rate v.s. number of users to average per round (source model trained in centralized manner with 30 client’s data).
(b) ResNet50: transfer rate v.s. maximum number of classes per user (10 users)..
(c) ResNet50: transfer rate v.s. the maximum number of classes per user (100 users).
(d) CNN: transfer rate v.s. the number of total users in the partition.
(e) CNN: transfer rate v.s. dirichlet alpha.
(f) CNN: transfer rate v.s. unbalance sgm.
Figure 9: Linear regression visualization
(a) CNN: transfer rate v.s. number of users to average per round (source model trained in centralized manner with full training dataset).
(b) CNN: transfer rate v.s. number of users to average per round (source model trained in centralized manner with 30 client’s data).
(c) CNN: transfer rate v.s. maximum number of classes per user (10 users)..
(d) CNN: transfer rate v.s. the maximum number of classes per user (100 users).
Figure 10: Linear regression visualization

### v2

@@ Appendix C Statistical Hypothesis Testing on Spearman correlation coefficient
$\mathbb{E}_{x^{\prime},{y}}$ $\mathbb{E}_{\mathcal{T}}[\|y-f_{n}(x^{\prime})\|^{2}_{2}]=\underbrace{\mathbb{E}_{x^{\prime},{y}}[\|y-\bar{f_{n}}(x^{\prime})\|^{2}_{2}]}_{\text{Bias}^{2}}$ (20)
$+\underbrace{\mathbb{E}_{x^{\prime},\mathcal{T}}[\|\bar{f}_{n}(x^{\prime})-f_{n}(x^{\prime})\|^{2}_{2}]}_{\text{Variance}}$
$=\mathbb{E}_{x^{\prime},{y}}[\|y-\bar{f}(x^{\prime})\|^{2}_{2}]+\frac{1}{n}\mathbb{E}_{x^{\prime},\mathcal{T}}[\|\bar{f}(x^{\prime})-f(x^{\prime},\mathcal{T})\|^{2}_{2}]$
Inequality 6 holds due to that $\text{Var}(\frac{1}{n}\sum_{i=1}^{n}f_{i}(\cdot;\mathcal{T}_{i}))=\frac{1}{n^{2}}\text{Var}(\sum_{i=1}^{n}f_{i}(\cdot;\mathcal{T}_{i}))$ and since each $f_{i}$ is trained independently, $\frac{1}{n^{2}}\text{Var}(\sum_{i=1}^{n}f_{i}(\cdot;\mathcal{T}_{i}))=\frac{1}{n^{2}}\sum_{i=1}^{n}\text{Var}(f(\cdot;\mathcal{T}))=\frac{1}{n}\text{Var}(f(\cdot;\mathcal{T}))$. ∎
Appendix F Attacking Communication-efficient FL Methods
We also explore whether our attack performs well on communication-efficient federated methods such as DAdaQuant Hönig et al. (2022) and FedPAQ Reisizadeh et al. (2020b). Specifically, we compare the standard FedAvg, Krum, Trimmed Mean and FedPAQ Reisizadeh et al. (2020b).
Figure 9: Attacking communication-efficient FL.
Appendix G Comparison between transfer attack and attack using traces of training.
In this section, we explore using the traces of training (checkpoints during the training of the federated model) to attack and compare with the transfer attack presented in the paper. We must emphasize that the former is a different threat model with the capability of keeping traces of the model at different stages of training. That is to say, the attack is a purely white-box one.
We present the attack result using checkpoints of different stages of the training as shown in Tab 4. We early stopped the model when the accuracy reaches 90% which is around 75 epochs. Then we use the checkpoints from 20, 40 and 60 epochs to perform the attack.
white-box 20e 40e 60e
transfer rate 91.48 89.98 91.45 94.66
1% transfer 10% transfer 30% transfer 100% transfer
transfer rate 2.49 57.69 83.91 68.94
Table 4: Comparison between using checkpoint (traces of training) attack and transfer attack.
Appendix H linear regression to visualize the correlation
@@ Appendix F Attacking Communication-efficient FL Methods
Figure 11: ResNet50: transfer rate v.s. number of users to average per round (source model trained in centralized manner with 30 client’s data).
Figure 12: ResNet50: transfer rate v.s. maximum number of classes per user (10 users).
Figure 13: ResNet50: transfer rate v.s. the maximum number of classes per user (100 users).
Figure 14: CNN: transfer rate v.s. the number of total users in the partition.
Figure 15: CNN: transfer rate v.s. dirichlet alpha.
Figure 16: CNN: transfer rate v.s. unbalance sgm.
Figure 17: CNN: transfer rate v.s. number of users to average per round (source model trained in centralized manner with full training dataset).
Figure 18: CNN: transfer rate v.s. number of users to average per round (source model trained in centralized manner with 30 client’s data).
Figure 19: CNN: transfer rate v.s. maximum number of classes per user (10 users).
Figure 20: CNN: transfer rate v.s. the maximum number of classes per user (100 users).